OpenStudy (anonymous):

approximation of 1/fourth root of 15

OpenStudy (anonymous):

let y = 15^1/4 ... take log on both sides

OpenStudy (mertsj):

Are you supposed to calculate this or solve by making successive guesses?

OpenStudy (anonymous):

calculate i suppose...

OpenStudy (amistre64):

\[\frac{15}{4}=\frac{12+3}{4}=\frac{12}{4}+\frac34\]

OpenStudy (amistre64):

lol, quarter ROOT ..... got it

OpenStudy (anonymous):

^ its the 1/4th root .. not 1/4 times...

OpenStudy (amistre64):

i like the 1/4 times alot better tho